Hi,
I've been out of school for 2 years and I cannot remember how to do the cube root of numbers.
The problem I have asks what is 4 times the cube root of 81.

Answer by Edwin McCravy(20054)

  ___      ___     _______     __________    
4∛81  = 4∛9*9 = 4∛3*3*3*3 = 4∛(3*3*3)*3 = 
  ____       __  _        __      _  
4∛3³*3 = 4*∛3³*∛3 = 4*3*∛3 = 12∛3

And you leave it like that.

Edwin